OpenStudy (anonymous):

Find an equation of the ellipse having the given points as foci and the given sum of the focal radii (-9, 0); (9, 0); 30

OpenStudy (amistre64):



OpenStudy (amistre64):

can you figure out the parts of: a,b,c from this drawing?

OpenStudy (amistre64):

since the midpoint between the foci is the center, is 0,0 the equation relates to \[\frac{x^2}{a^2}+\frac{y^2}{b^2}=1\] c = distance from center to a focus

OpenStudy (anonymous):

thanks :)